OpenStudy (anonymous):

rectangle as 5inches by 8inches area is?

jimthompson5910 (jim_thompson5910):

For a rectangle Area = Length * Width

jimthompson5910 (jim_thompson5910):

Area = Length * Width A = L*W A = 5*8 A = 40 So the area is 40 square inches. Don't forget to mention the units of "square inches" for the area.